OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

-7(x+2)=y-2 y=-7x-12

OpenStudy (amistre64):

... standard form is not slope intercept form